### v1

@@ 8 Rank robustness against perturbations
Let $A\in\mathbb{C}^{n\times n}$ be cramped with phases in $(-\pi,\pi)$. Then $\mathrm{rank}\,(I+AB)=n$ for all $B\in\mathcal{B}[\gamma]\cup\mathcal{C}_{1}[\alpha]$ with $\gamma>0$, $\alpha\in[0,\pi)$, if and only if $\gamma<{\sigma_{1}(A)}^{-1}$ and $\alpha<\min\{\pi-\phi_{1}(A),\pi+\phi_{n}(A)\}$.
@@ 9 Kronecker and Hadamard products
It is known that the singular values of $A\otimes B$ are given by $\sigma_{i}(A)\sigma_{j}(B),1\leq i\leq n,1\leq j\leq m$ [13, Theorem 4.2.15]. Regarding the phases of $A\otimes B$, we have the following result.
Theorem 9.1.
Let $A\in\mathbb{C}^{n\times n}$ and $B\in\mathbb{C}^{m\times m}$ be cramped matrices with phases in $(\theta_{1},\theta_{1}+\pi)$ and $(\theta_{2},\theta_{2}+\pi)$, where $\theta_{1}\in[-\pi,\delta(A))$ and $\theta_{2}\in[-\pi,\delta(B))$, respectively. If $\phi_{1}(A)+\phi_{1}(B)-\phi_{n}(A)-\phi_{n}(B)<\pi$, then $A\otimes B$ is cramped and its phases are given by $\phi_{i}(A)+\phi_{j}(B),1\leq i\leq n,1\leq j\leq m$.
Proof 9.2.
Let $A=T^{*}DT$ and $B=R^{*}ER$ be sectoral decompositions of $A$ and $B$, respectively. Then,
(20) $A\otimes B=(T^{*}DT)\otimes(R^{*}ER)=(T\otimes R)^{*}(D\otimes E)(T\otimes R),$
@@
Since $\phi_{1}(A)+\phi_{1}(B)-\phi_{n}(A)-\phi_{n}(B)<\pi$, it follows that $A\otimes B$ is cramped and (20) is a sectoral decomposition of $A\otimes B$. Furthermore, the phases of $A\otimes B$ are given by $\angle(\lambda_{i}(D)\lambda_{j}(E))$, i.e., $\phi_{i}(A)+\phi_{j}(B)$, $1\leq i\leq n,1\leq j\leq m$.
As for the Hadamard product $A\odot B$, a notably elegant result on its singular values is that [13, Theorem 5.5.4]
@@
Theorem 9.3.
Let $A,B\in\mathbb{C}^{n\times n}$ be cramped matrices with phases in $(\theta_{1},\theta_{1}+\pi)$ and $(\theta_{2},\theta_{2}+\pi)$, where $\theta_{1}\in[-\pi,\delta(A))$ and $\theta_{2}\in[-\pi,\delta(B))$, respectively. If $\phi_{1}(A)+\phi_{1}(B)-\phi_{n}(A)-\phi_{n}(B)<\pi$, then $A\odot B$ is cramped,
@@
Proof 9.4.
The claim follows from Theorems 9.1 and 4.1, since $A\odot B$ can be expressed as a compression of $A\otimes B$. We offer an alternative proof below.
Let $x\in\mathbb{C}^{n},x\neq 0$. By [13, Lemma 5.1.5],
@@ 10 Conclusions
In this paper, we define the phases of a cramped matrix and study their properties. We introduce the symmetric polar decomposition and generalized Cholesky factorization of a cramped matrix and establish their uniqueness. The symmetric polar decomposition seems to have advantages over the usual polar decomposition, at least in defining the matrix phases. In the scalar case, the symmetric polar decomposition takes the form $c=\sqrt{\sigma}e^{i\phi}\sqrt{\sigma}$. A number of useful properties of the matrix phases have been studied, including those of compressions, Schur complements, matrix products and sums. The rank robustness of a matrix against magnitude/phase perturbations, motivated from applications in robust control, has also been examined.
@@
The definition of phases can be generalized to matrices whose numerical ranges contain the origin on their boundaries. We call such matrices semicramped. A number of results in this paper have potential extensions to semicramped matrices, including those on sectoral and symmetric polar decompositions, matrix products, sums, and rank robustness.
@@
The definition of phases can be further extended to some matrices which are not cramped, such as block diagonal matrices with cramped diagonal blocks, the Kronecker products of cramped matrices, compound cramped matrices, etc. These matrices are constructed from cramped matrices and their phases can be defined by exploiting the phases of the original cramped matrices.
@@
How to define phases for general noncramped unitoid matrices, which are congruent to unitary matrices, remains open. A critical issue in this problem involves determining how the phases take values and deriving their corresponding properties. We expect that the notion of Riemann surface would play an important role in studying the phases of unitoid matrices.
Acknowledgements
@@
The authors wish to thank Axel Ringh and Chi-Kwong Li for valuable discussions.
?refname?
  • [1] B. D. Anderson and S. Vongpanitlerd, Network Analysis and Synthesis: A Modern Systems Theory Approach, Englewood Cliffs, NJ, USA: Prentice-Hall, Inc., 1973.
  • [2] S. K. Berberian, The numerical range of a normal operator, Duke Mathematical Journal, 31 (1964), pp. 479–483.
  • [3] R. Bhatia, Matrix Analysis, Springer-Verlag, New York, 1997.
  • [4] R. Bhatia, Linear algebra to quantum cohomology: The story of Alfred Horn’s inequalities, The American Mathematical Monthly, 108 (2001), pp. 289–318.
  • [5] C. DePrima and C. R. Johnson, The range of $A^{-1}A^{*}$ in $GL(n,C)$, Linear Algebra and its Applications, 9 (1974), pp. 209–222.
  • [6] W. A. Donnell, Minimax principles of the arguments of the proper values of a normal linear transformation, PhD thesis, Texas Tech University, 1971.
  • [7] S. Drury, Fischer determinantal inequalities and Higham’s conjecture, Linear Algebra and its Applications, 439 (2013), pp. 3129–3133.
  • [8] M. Fiedler, Special Matrices and Their Applications in Numerical Mathematics, 2nd ed., Dover Publications, Inc., 2008.
  • [9] S. Furtado and C. R. Johnson, Spectral variation under congruence, Linear and Multilinear Algebra, 49 (2001), pp. 243–259.
  • [10] M. Green and D. J. Limebeer, Linear Robust Control, Englewood Cliffs, NJ, USA: Prentice-Hall, Inc., 1995.
  • [11] A. Horn and R. Steinberg, Eigenvalues of the unitary part of a matrix, Pacific Journal of Mathematics, 9 (1959), pp. 541–550.
  • [12] R. A. Horn and C. R. Johnson, Matrix Analysis, Cambridge University Press, 1990.
  • [13] R. A. Horn and C. R. Johnson, Topics in Matrix Analysis, Cambridge University Press, 1991.
  • [14] C. R. Johnson and S. Furtado, A generalization of Sylvester’s law of inertia, Linear Algebra and its Applications, 338 (2001), pp. 287–290.
  • [15] M. Marcus and R. Thompson, On a classical commutator result, Journal of Mathematics and Mechanics, 16 (1966), pp. 583–588.
  • [16] A. W. Marshall, I. Olkin, and B. C. Arnold, Inequalities: Theory of Majorization and Its Applications, Springer, 1979.
  • [17] A. A. Nudelman and P. Shvartsman, The spectrum of the product of unitary matrices, Uspekhi Matematicheskikh Nauk, 13 (1958), pp. 111–117.
  • [18] I. Postlethwaite, J. Edmunds, and A. MacFarlane, Principal gains and principal phases in the analysis of linear multivariable feedback systems, IEEE Transactions on Automatic Control, 26 (1981), pp. 32–46.
  • [19] G. Stewart and J.-G. Sun, Matrix Perturbation Theory, Academic Press, 1990.
  • [20] O. Taussky, Commutators of unitary matrices which commute with one factor, Journal of Mathematics and Mechanics, (1961), pp. 175–178.
  • [21] R. Thompson, On the eigenvalues of a product of unitary matrices I, Linear and Multilinear Algebra, 2 (1974), pp. 13–24.
  • [22] F. Zhang, The Schur Complement and Its Applications, Springer Science & Business Media, 2006.
  • [23] F. Zhang, A matrix decomposition and its applications, Linear and Multilinear Algebra, 63 (2015), pp. 2033–2042.
  • [24] K. Zhou, J. C. Doyle, and K. Glover, Robust and Optimal Control, Upper Saddle River, NJ, USA: Prentice-Hall, Inc., 1996.

### v2

@@ 8 Rank robustness against perturbations
Let $A\in\mathbb{C}^{n\times n}$ be sectorial with phases in $(-\pi,\pi)$. Then $\mathrm{rank}\,(I+AB)=n$ for all $B\in\mathcal{B}[\gamma]\cup\mathcal{C}_{1}[\alpha]$ with $\gamma>0$, $\alpha\in[0,\pi)$, if and only if $\gamma<{\sigma_{1}(A)}^{-1}$ and $\alpha<\min\{\pi-\phi_{1}(A),\pi+\phi_{n}(A)\}$.
9 Sectorial matrix completion
The matrix completion problem is to determine the unspecified entries of a partial matrix so that the completed matrix has certain desired properties. Consider partial matrices of the form
$C=\left[\begin{matrix}C_{11}&\cdots&C_{1,1+p}&&?\\ \vdots&&&\ddots&\\ C_{p+1,1}&&&&C_{n-p,n}\\ &\ddots&&&\vdots\\ ?&&C_{n,n-p}&\cdots&C_{nn}\end{matrix}\right],$
where $C_{ij}\in\mathbb{C}^{n_{i}\times n_{j}}$ is specified for $|i-j|\leq p$ with $p$ a fixed integer $0\leq p\leq n-1$. The unspecified blocks are represented by question marks. Such partial matrices are called $p$-banded. See [27] for comprehensive discussions of matrix completions. It is shown in [28] that $C$ admits a positive semidefinite completion if and only if
$\left[\begin{matrix}C_{ll}&\cdots&C_{l,l+p}\\ \vdots&&\vdots\\ C_{l+p,l}&\cdots&C_{l+p,l+p}\end{matrix}\right]\geq 0,\text{ for }l=1,\dots,n-p.$
In a similar spirit, we have the following result.
Theorem 9.1
A $p$-banded partial matrix $C$ admits a completion in $\mathcal{C}[\alpha,\beta]$ with $0\leq\beta-\alpha<\pi$ if and only if
$\left[\begin{matrix}C_{ll}&\cdots&C_{l,l+p}\\ \vdots&&\vdots\\ C_{l+p,l}&\cdots&C_{l+p,l+p}\end{matrix}\right]\in\mathcal{C}[\alpha,\beta],\text{ for }l=1,\dots,n-p.$
Proof 10
The necessity follows from Lemma 4.1. It remains to show the sufficiency. Without loss of generality, assume $n=3$ and $p=1$. The general case can be shown by an additional induction.
Let
$C=\left[\begin{matrix}C_{11}&C_{12}&X_{13}\\ C_{21}&C_{22}&C_{23}\\ Y_{31}&C_{32}&C_{33}\end{matrix}\right],$
where $X_{13}$ and $Y_{31}$ are to be determined so that $C\in\mathcal{C}[\alpha,\beta]$, which is equivalent to requiring the two inequalities
$e^{i(\frac{\pi}{2}-\beta)}C+e^{-i(\frac{\pi}{2}-\beta)}C^{*}$ $\geq 0,$ (21)
$e^{i(-\frac{\pi}{2}-\alpha)}C+e^{-i(-\frac{\pi}{2}-\alpha)}C^{*}$ $\geq 0,$ (22)
hold simultaneously. To find such $X_{13}$ and $Y_{31}$, we partition $e^{i(\frac{\pi}{2}-\beta)}C+e^{-i(\frac{\pi}{2}-\beta)}C^{*}$ with compatible dimensions into
$e^{i(\frac{\pi}{2}-\beta)}C+e^{-i(\frac{\pi}{2}-\beta)}C^{*}=\left[\begin{matrix}A&B&X\\ B^{*}&E&F\\ X^{*}&F^{*}&G\end{matrix}\right].$
When $\begin{bmatrix}C_{11}&C_{12}\\ C_{21}&C_{22}\end{bmatrix},\begin{bmatrix}C_{22}&C_{23}\\ C_{32}&C_{33}\end{bmatrix}\in\mathcal{C}[\alpha,\beta]$, both $\begin{bmatrix}A&B\\ B^{*}&E\end{bmatrix}$ and $\begin{bmatrix}E&F\\ F^{*}&G\end{bmatrix}$ are positive semidefinite. By [29, Theorem 3.2], if we let $X=BE^{\dagger}F$, i.e.,
$e^{i(\frac{\pi}{2}-\beta)}X_{13}+e^{-i(\frac{\pi}{2}-\beta)}Y^{*}_{31}=BE^{\dagger}F,$ (23)
where $E^{\dagger}$ is the Moore-Penrose pseudoinverse of $E$, then the inequality (21) holds. Similarly, we partition
$e^{i(-\frac{\pi}{2}-\alpha)}C+e^{-i(-\frac{\pi}{2}-\alpha)}C^{*}=\left[\begin{matrix}\tilde{A}&\tilde{B}&\tilde{X}\\ \tilde{B}^{*}&\tilde{E}&\tilde{F}\\ \tilde{X}^{*}&\tilde{F}^{*}&\tilde{G}\end{matrix}\right$]
and let $\tilde{X}=\tilde{B}\tilde{E}^{\dagger}\tilde{F}$, i.e.,
$e^{i(-\frac{\pi}{2}-\alpha)}X_{13}+e^{-i(-\frac{\pi}{2}-\alpha)}Y^{*}_{31}=\tilde{B}\tilde{E}^{\dagger}\tilde{F}$. (24)
Then the inequality (22) holds. Finally, solving equations (23) and (24) together yields desired $X_{13}$ and $Y_{31}$ so that $C\in\mathcal{C}[\alpha,\beta]$.
10 Sectorial matrix decomposition
In this section, we discuss a matrix decomposition problem, which can be regarded as a dual problem of the sectorial matrix completion studied in the previous section. A $p$-banded matrix
$C=\left[\begin{matrix}C_{11}&\cdots&C_{1,p+1}&&0\\ \vdots&&&\ddots&\\ C_{p+1,1}&&&&C_{n-p,n}\\ &\ddots&&&\vdots\\ 0&&C_{n,n-p}&\cdots&C_{nn}\end{matrix}\right],$
where $C_{ij}\in\mathbb{C}^{n_{i}\times n_{j}}$, is said to admit a positive semidefinite decomposition if it holds $C=C_{1}+\dots+C_{n-p}$ as in Figure 2, with $C_{l}=\mathrm{diag}\{0,\tilde{C}_{l},0\}$ and $\tilde{C}_{l}\geq 0$ for $l=1,\dots,n-p$. It is shown in [30] that $C$ admits a positive semidefinite decomposition if and only if $C\geq 0$.
Figure 2: Banded matrix decomposition.
Extending to the realm of sectorial matrices, we say a matrix $C$ admits a decomposition in $\mathcal{C}[\alpha,\beta]$ with $0\leq\beta-\alpha<\pi$ if it holds $C=C_{1}+\dots+C_{n-p}$ with $C_{l}=\mathrm{diag}\{0,\tilde{C}_{l},0\}$ and $\tilde{C}_{l}\in\mathcal{C}[\alpha,\beta]$ for $l=1,\dots,n-p$. We have the following result.
Theorem 10.1
A $p$-banded matrix $C$ admits a decomposition in $\mathcal{C}[\alpha,\beta]$ if and only if $C\in\mathcal{C}[\alpha,\beta]$.
Proof 11
We first show the necessity. From $\tilde{C}_{l}\in\mathcal{C}[\alpha,\beta]$, we have
$e^{i(\frac{\pi}{2}-\beta)}\tilde{C}_{l}+e^{-i(\frac{\pi}{2}-\beta)}\tilde{C}^{*}_{l}\geq 0\text{ and }e^{i(-\frac{\pi}{2}-\alpha)}\tilde{C}_{l}+e^{-i(-\frac{\pi}{2}-\alpha)}\tilde{C}^{*}_{l}\geq 0.$
Then,
$\begin{split}e^{i(\frac{\pi}{2}-\beta)}C+e^{-i(\frac{\pi}{2}-\beta)}C^{*}&=\sum_{l=1}^{n-p}\left(e^{i(\frac{\pi}{2}-\beta)}C_{l}+e^{-i(\frac{\pi}{2}-\beta)}C_{l}^{*}\right)\geq 0,\\ e^{i(-\frac{\pi}{2}-\alpha)}C+e^{-i(-\frac{\pi}{2}-\alpha)}C^{*}&=\sum_{l=1}^{n-p}\left(e^{i(-\frac{\pi}{2}-\alpha)}C_{l}+e^{-i(-\frac{\pi}{2}-\alpha)}C_{l}^{*}\right)\geq 0,\end{split}$
from which $C\in\mathcal{C}[\alpha,\beta]$ follows.
We will show the sufficiency by construction. Without loss of generality, assume $n=3$ and $p=1$. The general case can be shown by an additional induction.
Let
$C=\left[\begin{matrix}C_{11}&C_{12}&0\\ C_{21}&C_{22}&C_{23}\\ 0&C_{32}&C_{33}\end{matrix}\right]=\left[\begin{matrix}C_{11}&C_{12}&0\\ C_{21}&X_{22}&0\\ 0&0&0\end{matrix}\right]+\left[\begin{matrix}0&0&0\\ 0&Y_{22}&C_{23}\\ 0&C_{32}&C_{33}\end{matrix}\right],$
where $X_{22}$ and $Y_{22}$ are to be determined so that $\begin{bmatrix}C_{11}&C_{12}\\ C_{21}&X_{22}\end{bmatrix},\begin{bmatrix}Y_{22}&C_{23}\\ C_{32}&C_{33}\end{bmatrix}\in\mathcal{C}[\alpha,\beta]$. By $C\in\mathcal{C}[\alpha,\beta]$, we have
$e^{i(\frac{\pi}{2}-\beta)}C+e^{-i(\frac{\pi}{2}-\beta)}C^{*}$ $\geq 0,$ (25)
$e^{i(-\frac{\pi}{2}-\alpha)}C+e^{-i(-\frac{\pi}{2}-\alpha)}C^{*}$ $\geq 0.$ (26)
Partitioning $e^{i(\frac{\pi}{2}-\beta)}C+e^{-i(\frac{\pi}{2}-\beta)}C^{*}$ with compatible dimensions into
$e^{i(\frac{\pi}{2}-\beta)}C+e^{-i(\frac{\pi}{2}-\beta)}C^{*}=\left[\begin{matrix}A&B&0\\ B^{*}&E&F\\ 0&F^{*}&G\end{matrix}\right]=\left[\begin{matrix}A&B&0\\ B^{*}&X&0\\ 0&0&0\end{matrix}\right]+\left[\begin{matrix}0&0&0\\ 0&Y&F\\ 0&F^{*}&G\end{matrix}\right],$
we have $\begin{bmatrix}A&B\\ B^{*}&E\end{bmatrix}\geq 0$, which implies $\mathcal{R}(B)\subset\mathcal{R}(A)$ [17, Theorem 1.19], where $\mathcal{R}$ denotes the range. Moreover, if we set $X=B^{*}A^{\dagger}B$, we can see that the generalized Schur complement of $A$ in $\begin{bmatrix}A&B\\ B^{*}&X\end{bmatrix}$ equals 0. Thus, by a property of generalized Schur complement [17, Theorem 1.20], we have $\begin{bmatrix}A&B\\ B^{*}&X\end{bmatrix}\geq 0$.
It remains to show that $Y\geq 0$ and $\begin{bmatrix}Y&F\\ F^{*}&G\end{bmatrix}\geq 0$. Since $Y=E-X=E-B^{*}A^{\dagger}B$ is the generalized Schur complement of $A$ in $\begin{bmatrix}A&B\\ B^{*}&E\end{bmatrix}$, we have $Y\geq 0$. Furthermore, by
$\left[\begin{matrix}A&B&0\\ B^{*}&E&F\\ 0&F^{*}&G\end{matrix}\right]\geq 0,$
we can see
$\begin{bmatrix}E&F\\ F^{*}&G\end{bmatrix}-\begin{bmatrix}B^{*}\\ 0\end{bmatrix}A^{\dagger}\begin{bmatrix}B&0\end{bmatrix}=\begin{bmatrix}E-B^{*}A^{\dagger}B&F\\ F^{*}&G\end{bmatrix}=\begin{bmatrix}Y&F\\ F^{*}&G\end{bmatrix}\geq 0.$
Therefore, we have an equation of $X_{22}$:
$e^{i(\frac{\pi}{2}-\beta)}X_{22}+e^{-i(\frac{\pi}{2}-\beta)}X_{22}^{*}=B^{*}A^{\dagger}B.$ (27)
Similarly, we partition
$e^{i(-\frac{\pi}{2}-\alpha)}C+e^{-i(-\frac{\pi}{2}-\alpha)}C^{*}=\left[\begin{matrix}\tilde{A}&\tilde{B}&0\\ \tilde{B}^{*}&\tilde{E}&\tilde{F}\\ 0&\tilde{F}^{*}&\tilde{G}\end{matrix}\right]=\left[\begin{matrix}\tilde{A}&\tilde{B}&0\\ \tilde{B}^{*}&\tilde{X}&0\\ 0&0&0\end{matrix}\right]+\left[\begin{matrix}0&0&0\\ 0&\tilde{Y}&\tilde{F}\\ 0&\tilde{F}^{*}&\tilde{G}\end{matrix}\right$]
and let $\tilde{X}=\tilde{B}^{*}\tilde{A}^{\dagger}\tilde{B}$, i.e.,
$e^{i(-\frac{\pi}{2}-\alpha)}X_{22}+e^{-i(-\frac{\pi}{2}-\alpha)}X_{22}^{*}=\tilde{B}^{*}\tilde{A}^{\dagger}\tilde{B}$. (28)
Then $\begin{bmatrix}\tilde{A}&\tilde{B}\\ \tilde{B}^{*}&\tilde{X}\end{bmatrix}\geq 0,\begin{bmatrix}\tilde{Y}&\tilde{F}\\ \tilde{F}^{*}&\tilde{G}\end{bmatrix}\geq 0$. Finally, solving equations (27) and (28) together yields desired $X_{22}$.
@@ 11 Kronecker and Hadamard products
It is known that the singular values of $A\otimes B$ are given by $\sigma_{i}(A)\sigma_{j}(B),1\leq i\leq n,1\leq j\leq m$ [6, Theorem 4.2.15]. Regarding the phases of $A\otimes B$, we have the following result.
Theorem 11.1
Let $A\in\mathbb{C}^{n\times n}$ and $B\in\mathbb{C}^{m\times m}$ be sectorial matrices with phases in $(\theta_{1},\theta_{1}+\pi)$ and $(\theta_{2},\theta_{2}+\pi)$, where $\theta_{1}\in[-\pi,\delta(A))$ and $\theta_{2}\in[-\pi,\delta(B))$, respectively. If $\phi_{1}(A)+\phi_{1}(B)-\phi_{n}(A)-\phi_{n}(B)<\pi$, then $A\otimes B$ is sectorial and its phases are given by $\phi_{i}(A)+\phi_{j}(B),1\leq i\leq n,1\leq j\leq m$.
Proof 12
Let $A=T^{*}DT$ and $B=R^{*}ER$ be sectorial decompositions of $A$ and $B$, respectively. Then,
$A\otimes B=(T^{*}DT)\otimes(R^{*}ER)=(T\otimes R)^{*}(D\otimes E)(T\otimes R),$ (29)
@@
Since $\phi_{1}(A)+\phi_{1}(B)-\phi_{n}(A)-\phi_{n}(B)<\pi$, it follows that $A\otimes B$ is sectorial and (29) is a sectorial decomposition of $A\otimes B$. Furthermore, the phases of $A\otimes B$ are given by $\angle(\lambda_{i}(D)\lambda_{j}(E))$, i.e., $\phi_{i}(A)+\phi_{j}(B)$, $1\leq i\leq n,1\leq j\leq m$.
As for the Hadamard product $A\odot B$, a notably elegant result on its singular values is that [6, Theorem 5.5.4]
@@
Theorem 11.3
Let $A,B\in\mathbb{C}^{n\times n}$ be sectorial matrices with phases in $(\theta_{1},\theta_{1}+\pi)$ and $(\theta_{2},\theta_{2}+\pi)$, where $\theta_{1}\in[-\pi,\delta(A))$ and $\theta_{2}\in[-\pi,\delta(B))$, respectively. If $\phi_{1}(A)+\phi_{1}(B)-\phi_{n}(A)-\phi_{n}(B)<\pi$, then $A\odot B$ is sectorial,
@@
Proof 13
The claim follows from Theorems 11.1 and Lemma 4.1, since $A\odot B$ can be expressed as a compression of $A\otimes B$. We offer an alternative proof below.
Let $x\in\mathbb{C}^{n},x\neq 0$. By [6, Lemma 5.1.5],
@@ 12 Conclusions
In this paper, we define the phases of a sectorial matrix and study their properties. We introduce the symmetric polar decomposition and generalized Cholesky factorization of a sectorial matrix and establish their uniqueness. The symmetric polar decomposition seems to have advantages over the usual polar decomposition, at least in defining the matrix phases. In the scalar case, the symmetric polar decomposition takes the form $c=\sqrt{\sigma}e^{i\phi}\sqrt{\sigma}$. A number of useful properties of the matrix phases have been studied, including those of compressions, Schur complements, matrix products and sums. The rank robustness of a matrix against magnitude/phase perturbations, motivated from applications in robust control, has also been examined. A pair of problems: sectorial matrix completion and decomposition extending those of the positive semidefinite completion and decomposition are studied.
@@
The definition of phases can be generalized to matrices whose numerical ranges contain the origin on their boundaries. We call such matrices semi-sectorial. A number of results in this paper have potential extensions to semi-sectorial matrices, including those on sectorial and symmetric polar decompositions, matrix products, sums, and rank robustness.
@@
The definition of phases can be further extended to some matrices which are not sectorial, such as block diagonal matrices with sectorial diagonal blocks, the Kronecker products of sectorial matrices, compound sectorial matrices, etc. These matrices are constructed from sectorial matrices and their phases can be defined by exploiting the phases of the original sectorial matrices.
@@
How to define phases for general nonsectorial unitoid matrices, which are congruent to unitary matrices, remains open. A critical issue in this problem involves determining how the phases take values and deriving their corresponding properties. We expect that the notion of Riemann surface would play an important role in studying the phases of unitoid matrices.
Acknowledgment
This work was partially supported by the Research Grants Council of Hong Kong Special Administrative Region, China, under the Theme-Based Research Scheme T23-701/14-N and the General Research Fund 16211516.
@@ Acknowledgment
The authors also wish to thank Di Zhao, Axel Ringh, and Chi-Kwong Li for valuable discussions.